OpenStudy (anonymous):

PLEASE HELP!!! REALLY QUICK! What are the exact solutions of x2 - x - 4 = 0?

OpenStudy (anonymous):

I believe D is the correct answer

OpenStudy (anonymous):

Well, do you remember what the formula for finding the two solutions of a quadratic equation looked like ?

OpenStudy (anonymous):

Given a general quadratic form, say: a*x^2 + b*x + c = 0, where a,b,c are real constants.

OpenStudy (anonymous):

yes it is\[x = -b \sqrt{b^2 - 4ac} \over 2a\]

OpenStudy (anonymous):

I suggest that - at least for the first 50-100 quadratics you will solve - to write down what a,b and c are explicitly because it's very easy to get tricked with them.

OpenStudy (anonymous):

in this equation a = 1 b= 1 c = 4

OpenStudy (anonymous):

Right, there's a +/- sight instead of multiplication there but I think it's just a typo. Very good though! Right, let's take a look at this equation

OpenStudy (anonymous):

oh yea ax^2 + bx +c

OpenStudy (anonymous):

Yea it's a typo i meant to put it there =)

OpenStudy (anonymous):

You have x^2 - x - 4 =0.

OpenStudy (anonymous):

\[x = -1+ - \sqrt{1^2 - 4(4)(1)} \over 2(1)\]

OpenStudy (anonymous):

If b were to equal 1 here, how would x^2 - x - 4 be then different from x^2 + x - 4

OpenStudy (anonymous):

hmmm, u lost me =)

OpenStudy (anonymous):

The best way to not get confused by this in the beginning is to write them down like this : 1*x^2 + (-1)*x + (-1)*4 =0. Do you agree that this is the same thing as the original equation you started out with ?

OpenStudy (anonymous):

What I am trying to say is that b=-1.

OpenStudy (radar):

What is wrong with C?

OpenStudy (anonymous):

And c=-4.

OpenStudy (anonymous):

lol okay hold on lemme write this down and figure it out =)

OpenStudy (anonymous):

Take your time!

OpenStudy (anonymous):

oh okay i see your point, yea that makes sense =)

OpenStudy (radar):

Note that b = -1 therefore -b = 1

OpenStudy (anonymous):

But if you graph x^2 +x -4 and x^2 - x -4 it is two diff parabolas so i don't see how that could be the same thing =) if u see where i'm coming from =)

OpenStudy (anonymous):

@AngusV

OpenStudy (anonymous):

My prob is that when i worked out the radicand i got -15 every time and neither of those answers are -15 in the radicand =)

OpenStudy (anonymous):

Yes, of course they are not the same thing! What I was trying to say was that they are two different things and they both can't have b=1. x^2 - x - 4 which has b= -1 is different from x^2 +x - 4 which has b = 1. My point was to illustrate that since those two equations aren't equivalent, neither is the b equivalent. And since b is not equivalent, neither are those two. If it makes any sense.

OpenStudy (anonymous):

Well, let's look at the radicand thing. There you have the square toot of b^2 - 4*a*c.

OpenStudy (anonymous):

so i'm trying to figure out how the sign could possibly change to make it = 17 . . . . . . =)

OpenStudy (anonymous):

When you plug in the values you have that the radicand thing equals : square root of (-1)^2 - 4*1*(-4)

OpenStudy (anonymous):

Oh okay i get your non - equality point between those two equations =)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Well, you forget that c is -4. That's why it changes the thing to 17. When you solve that you get that the radicand thing equals square root of 1 - (-16) which is square root of 1 + 16 = square root of 17.

OpenStudy (anonymous):

All in all, the answer is C.

OpenStudy (anonymous):

Ahhhhh i see, okay thanks so much for the help i think i finally get this =)

OpenStudy (radar):

I agree that C is a good choice. I used the "completing the square method" which I find easier for simple quadratics than the Quadratic fromula.

OpenStudy (radar):

In fact the Quadratic Formula was derived using the "completing the square"

OpenStudy (anonymous):

lol okay thanks so much, =)